OpenStudy (anonymous):

plz help will give medal Which of these sequences is a geometric sequence? 1, 2, 4, 7, 11, 16, 22, … 2, 4, 8, 14, 22, 38, … 3, 6, 9, 12, 15, 18, 21, … 3, 9, 27, 81, 343, 729, …

OpenStudy (sidsiddhartha):

for a geometric sequence \[\frac{ second~term }{ first~term }=3rd~term/2nd~term=4th~term/3rd ~term=constant\]

OpenStudy (sidsiddhartha):

so which one is it,?

OpenStudy (sidsiddhartha):

can u tell me?

OpenStudy (anonymous):

i chose c but it wasnt correct lol i suck at math

OpenStudy (sidsiddhartha):

look at the fourth one \[9/3=27/9=81/27=343/81=...=3\] so for this one the common ration is constant which is 3, so this is a geometric sequence

OpenStudy (anonymous):

oh ok ok i get it :)